What is bigger 32 hours or 1 1/4 days?

Answer by jim_thompson5910(35256)

1 1/4 days = 1 + 1/4

1 1/4 days = 4/4 + 1/4

1 1/4 days = 5/4 days

1 1/4 days = 5/4 * (24) hours

1 1/4 days = 120/4 hours

1 1/4 days = 30 hours

So because 1 1/4 days = 30 hours, this means the longer length of time is 32 hours